For the reaction of sodium with oxygen to produce sodium oxide, the rate of disappearance of sodium is 1.2 x 10-3 M/sec. What is the rate of appearance of sodium oxide? The rate of disappearance of oxygen? 4 Na + O2   -> 2 Na2O
Answer #1

4 Na + O2   -> 2 Na2O

-1/4\Delta[Na]/\DeltaT    = -\Delta[O2]/\DeltaT      = 1/2 \Delta [Na2O]/\DeltaT

-1/4 *1.2*10^-3 = -\Delta[O2]/\DeltaT

3*10^-4         = - \Delta [O2]/\DeltaT'

The rate of disappearance of oxygen = 3*10^-4 M/sec

-1/4\Delta[Na]/\DeltaT   = 1/2 \Delta [Na2O]/\DeltaT

-1/4*1.2*10^-3    = 1/2 \Delta [Na2O]/\DeltaT

3*10^-4            = 1/2 \Delta [Na2O]/\DeltaT

6*10^-4           = \Delta [Na2O]/\DeltaT

the rate of appearance of sodium oxide is 6*10^-4M/sec